The line is perpendicular to the line whose equation is 4x - y = 6 and has the same y-intercept as this line.
Found 2 solutions by funmath, tutorcecilia:
Answer by funmath(2873)

I am assuming that you have to find the equation of the line that meets those requirements.
Put the line in slope intercept form (y=mx+b)
m=slope and (0,b)=y-intecept
A line that is perpendicular to this one would have a slope that in the negative reciprocal of this slope (-1/m)
4x-y=6
-4x+4x-y=-4x+6
0-y=-4x+6
-y=-4x+6
-1(-y)=-1(-4x+6)
y=4x-6
The slope (m) of this line is 4 the slope of a line that is perpendicular to this line (-1/m) is -1/4.
The y intercept (b=-6) is the same, so the answer is:
y=-1/4x-6

Answer by tutorcecilia(2144)

Line #1: 4x-y=6
y=4x-6 [Re-write in the slope-intercept form of (y=mx+b)]
m=4
b=-6
.
Line #2 Slope is the negative reciprocal of m=4
m=-1/4
b=-6 ["has the same y-intercept" (-6)]
y=mx+b [Use the slope-intercept formula]
y=-1/4x-6